# Designing AI-Resilient Assessments Using Interconnected Problems: A Theoretically Grounded and Empirically Validated Framework

### v1

@@ Abstract
The proliferation of generative AI has rendered traditional, modular assessments in computing education obsolete, creating a critical disconnect between academic training and industry practice. This paper presents a theoretically grounded framework for designing AI-resilient assessments, supported by formal proofs and empirical validation. We make three primary contributions. First, we establish two formal theorems: (1) assessments composed of interconnected problems, where outputs serve as inputs to subsequent stages, are inherently more AI-resilient than modular assessments due to multi-step reasoning and context limitations of large language models; and (2) semi-structured problems with deterministic success criteria provide more reliable measures of student competency than fully open-ended projects, which allow AI systems to default to familiar solution patterns. These findings challenge widespread recommendations from international policy organizations [unesco2025assessment] and institutional guidelines [cornell2024ai, wustl2024ai] that open-ended, authentic assessments are the primary solution to maintaining academic integrity in the AI era. Second, we validate these theorems through empirical analysis of four university data science courses ($N=138$), demonstrating a substantial AI inflation effect: students achieve near-perfect scores (9090100100%) on AI-assisted modular homework but drop approximately $30$ percentage points on proctored exams (Cohen’s $d=1.51$). Critically, interconnected projects maintain strong alignment with modular assessments ($r=0.954$, $p<.001$) while remaining AI-resilient, whereas proctored exams show only moderate correlation ($r=0.726$, $p<.001$), suggesting they assess different competencies. Third, we operationalize these findings into a practical design procedure that translates theoretical principles into implementable strategies, providing computing educators with a systematic, evidence-based approach for designing assessments that encourage deeper engagement, authentically reflect industry practice, and naturally resist trivial AI delegation.
@@ I Introduction
Since the release of generative AI tools like ChatGPT in late 2022, the rapid adoption of this technology has reshaped professional practice at a fundamental level [mckinsey2024state, ey2024work]. Tasks that once defined early-career technical work—drafting boilerplate code, writing simple SQL queries, or executing small-scale data analysis—are now completed instantly by AI tools, a shift confirmed by large-scale industry experiments [dellacqua2023navigating]. Yet, these same modular tasks continue to dominate academic assessments, creating a widening and unsustainable gap between university curricula and actual industry requirements [jin2025generative]. This misalignment raises a critical question: What is the educational value of practicing skills that industry has already automated? Students quickly recognize this disconnect, undermining their motivation and learning when assignments feel obsolete [dempere2023impact].
@@
While educators widely acknowledge this gap, proposed solutions often focus on how instructors can use AI as a teaching assistant—for intelligent tutoring, adaptive feedback, or automated grading—rather than on how assessment design itself must fundamentally change [ZawackiRichter2019, BondEtAl2024]. Such an approach overlooks the core challenge: the nature of meaningful technical work has shifted. Professionals now focus less on executing discrete tasks and more on orchestrating complex, multi-step workflows that integrate AI as a component, a form of higher-order work that current AI systems cannot replicate [brynjolfsson2023macroeconomics]. This shift demands a new pedagogical focus on AI-resilient assessments: assignments that cultivate genuine learning precisely because they cannot be fully delegated to an AI.
Our approach is grounded in the recognized limitations of large language models (LLMs). While proficient at single-step, modular problems, even state-of-the-art models like GPT-$5$ exhibit brittleness in tasks requiring sustained, multi-step reasoning [phan2025humanity, shojaee2025illusion]. Empirical benchmarks consistently show that LLMs struggle with long inference chains, logical negation, and discovering abstract structural patterns, often exploiting superficial token biases rather than engaging in genuine inference [parmar2024logicbench, gendron2024abstractreasoners, jiang2024tokenbias]. This limitation is not merely theoretical; domain-specific evaluations in fields like biostatistics confirm that LLMs fail on multi-stage procedures unless guided by iterative human feedback [ignjatovic2023biostat]. The boundary of AI capability, therefore, defines the frontier of essential human skill.
@@
This paper argues that the key to AI-resilient assessment lies in designing projects with a high degree of interconnectedness. Based on empirical observations from teaching four sessions of upper-level courses in data science and statistical learning, we make two central claims. First, we contend that complexity, defined by the interdependence among a project’s components, is the primary mechanism for fostering the integrative problem-solving skills that current AI tools lack and that industry now demands. Second, we challenge the widespread recommendation—articulated in international policy documents [unesco2025assessment] and institutional guidelines [cornell2024ai, wustl2024ai]—that open-ended, authentic assessments are the primary solution to maintaining academic integrity in the age of generative AI. While these recommendations rest on the assumption that AI ‘struggles with original thought and complex synthesis’ [unesco2025assessment], recent evidence demonstrates that LLMs’ limitations stem not from an inability to generate creative responses, but from fundamental architectural constraints in multi-step reasoning and context management [phan2025humanity, shojaee2025illusion].
@@ II Related Work
Our work is situated at the intersection of three distinct but related research areas: ($1$) the impact of generative AI on academic integrity and assessment, including recent evidence that generative AI can harm learning when used without appropriate guardrails [bastani2025generative, Cotton2023, Sullivan2023, Yan2024]; ($2$) the principles of authentic assessment and complex problem-solving in STEM education [Ullah2020, Csapo2017, Herde2016]; and ($3$) the cognitive science of task complexity [Chen2023]. While no prior work directly proposes a framework for AI-resilience based on interconnectedness, these adjacent fields establish the foundation and necessity for our contribution.
@@ II-A Generative AI, Assessment, and Academic Integrity
The release of ChatGPT precipitated a surge of academic literature focused on the challenges generative AI poses to traditional assessment and academic integrity. A significant body of this work documents the scope of the problem, highlighting the ease with which students can use AI to complete assignments, thereby undermining learning and creating an integrity crisis [Cotton2023, Sullivan2023, BinNashwan2023]. These studies effectively diagnose the issue, arguing that the availability of powerful AI necessitates a fundamental rethinking of assessment practices. However, the proposed solutions often remain high-level, advocating for a shift towards in-class assessments, oral examinations, or a greater emphasis on process over product, without offering a systematic, scalable framework for redesigning coursework itself [Yan2024].
Moreover, overly focusing on in-class assessments may test a completely different skill—test-taking ability rather than the intended learning outcomes [bastani2025generative]. Forgoing project-based assessment deprives students of the opportunity to practice skills that closely mimic real-world project work by industry professionals [Ullah2020, argus2022embedding]. We argue that adding guardrails through designing AI-resilient assessments is more beneficial, as it preserves authentic learning experiences while addressing integrity concerns.
@@ II-B Authentic Assessment and Complex Problem-Solving
Our approach builds on a long tradition of research into authentic assessment and complex problem-solving. Authentic assessment emphasizes mirroring the challenges and standards of professional practice, a principle well-established in engineering and computing education [Ullah2020]. Similarly, the field of complex problem-solving (CPS) focuses on assessing cognitive processes in dynamic, non-routine situations, which aligns with our goal of moving beyond static, modular tasks [Csapo2017, Herde2016]. While philosophically aligned, this body of work predates the generative AI era and thus does not explicitly address how to design assessments that are resilient to AI delegation. It provides the pedagogical why—the need for real-world problem-solving—but not the operational how in a world where AI can solve many seemingly complex but isolated tasks.
@@ II-C Task Complexity and Cognitive Load Theory
To ground our concept of interconnectedness, we draw on Cognitive Load Theory (CLT), particularly the concept of element interactivity. CLT posits that the intrinsic difficulty of a task is determined by the number of interacting elements that must be simultaneously processed in working memory [Chen2023]. A task with low element interactivity (e.g., solving a series of independent coding problems) can be broken down and solved piece-by-piece. In contrast, a task with high element interactivity (e.g., debugging a system where a change in one module has cascading effects on others) requires a holistic understanding of the entire system. We argue that current generative AI excels at tasks with low element interactivity but struggles with high-interactivity problems. Therefore, designing assessments with high element interactivity—what we term interconnectedness—is a theoretically grounded approach to creating AI-resilient learning experiences. Building on this foundation, we now formalize these principles through two theorems that establish the theoretical basis for AI-resilient assessment design.
@@ III Methodology: A Framework for AI-Resilient Assessment
Our methodology is grounded in a theoretical framework that we first establish through two theorems. We then proceed to validate these theorems through a multi-year empirical study.
@@ III-A Theoretical Framework
We begin by formally stating and proving two theorems that underpin our assessment design principles.
Theorem 1.
@@
Proof of Theorem 1.
Let an assessment be a sequence of problems $P=(p_{1},p_{2},...,p_{n})$. In a modular design, each $p_{i}$ is independent. In an interconnected design, the solution to $p_{i+1}$ is contingent upon the solution of $p_{i}$. A generative AI model must therefore solve the entire sequence $(p_{1},...,p_{i})$ to correctly solve $p_{i}$. This requirement increases the computational and reasoning load on the model in two primary ways. First, it necessitates multi-step, sequential reasoning, a known weakness of current large language models (LLMs) where performance degrades as the number of steps increases [Wei2022, Kojima2022]. Second, the cumulative context required to solve later problems in the sequence can exceed the model’s context window, leading to information loss and failure [Narayanan2023]. Because a modular design allows each problem to be solved in a separate, stateless context, it avoids these vulnerabilities. Thus, an interconnected design is inherently more resilient to trivial AI delegation and better reflects the complex, stateful workflows common in professional practice [Brynjolfsson2023]. ∎
Theorem 2.
@@
Proof of Theorem 2.
Large language models (LLMs) are fundamentally trained to minimize perplexity on their training distribution, which biases them toward generating sequences that closely match patterns encountered during pre-training [Huang2025Verbatim, Carlini2022Quantifying]. When presented with a fully open-ended problem (e.g., “analyze a dataset of your choice”), the unconstrained (open-ended problem) solution space allows the model to select a problem instance that lies within the model’s training distribution, where memorized patterns and dataset artifacts can be exploited as shortcuts [Du2022Shortcut]. This phenomenon, known as shortcut learning, occurs when models rely on spurious correlations rather than robust reasoning, significantly degrading generalizability to out-of-distribution tasks [Du2022Shortcut, Tang2023Lazy].
Recent work has demonstrated that LLMs exhibit lazy learning behavior, preferentially following the path of least resistance when afforded flexibility in problem selection [Tang2023Lazy, Sanyal2025Path]. Specifically, Tang et al. [Tang2023Lazy] show that larger models are more likely to exploit shortcuts during inference, while Sanyal et al. [Sanyal2025Path] demonstrate that policy optimization in LLMs “consistently follows the path of least resistance,” avoiding complex reasoning when simpler alternatives are available. In the context of open-ended assessments, this translates to students (with AI assistance) selecting problem variants with low perplexity—those that align closely with the model’s training data—thereby producing high-quality outputs that do not reflect genuine problem-solving ability on novel or challenging tasks.
Furthermore, research on verbatim memorization reveals that sequences with lower perplexity are significantly more likely to be memorized and reproduced by LLMs [Huang2025Verbatim]. Huang et al. demonstrate that “sequences that are better recognized by the language model, i.e., sequences with lower perplexity, are more likely to be verbatim memorized,” and that this memorization is “intertwined with the LM’s general capabilities” [Huang2025Verbatim]. Consequently, when students are permitted to choose their own datasets or problem instances, they (or the AI tools they employ) will gravitate toward familiar, well-represented problem types in the training corpus, where the model can leverage memorized solution templates rather than engaging in novel reasoning.
A semi-open-ended project addresses this limitation by constraining the solution space through two mechanisms: (1) specifying a particular dataset or problem instance, and (2) defining deterministic success criteria (e.g., “achieve $90\%$ F1-score on this held-out test set”). This design forces engagement with a potentially out-of-distribution task, where the model cannot simply retrieve memorized solutions or exploit dataset artifacts. The deterministic evaluation metric further prevents the model from optimizing for plausibility (a subjective, perplexity-driven objective) rather than correctness (an objective, task-specific objective). By narrowing the optimization landscape to a specific, well-defined problem, semi-open-ended assessments compel students to demonstrate competency on the assigned task, yielding a more reliable and robust measure of their problem-solving ability [Ullah2020].
In summary, the adversarial robustness of semi-open-ended assessments stems from their resistance to the fundamental optimization behavior of LLMs: the tendency to exploit shortcuts, memorize training data patterns, and follow paths of least resistance. By constraining the problem space and imposing deterministic evaluation, we force both students and AI tools to engage with the specific complexities of the task, providing a more authentic measure of competency in an AI-saturated educational environment. ∎
@@ III-B Empirical Validation
To validate these theorems, we analyzed assessment data from four offerings of upper-level undergraduate and graduate courses in data science spanning multiple academic years. This section details the course context, the assessment designs, and the data analysis procedures.
@@ III-B1 Study Context and Population
The study was conducted across four course offerings at a large research university between Fall 2024 and Spring 2025:
@@
  • Data Mining (Fall 2024): An undergraduate course (N=41) also with a mix of modular assignments and an open-ended final project.
@@
  • Predictive Analytics (Fall 2025): The same undergraduate course (N=22) but with a mix of modular assignments and a redesigned, interconnected final project.
III-B2 Data Collection and Analysis
For each course, we collected anonymized student grade data for all assessments. Our analysis proceeds in three stages:
@@ III-B2 Data Collection and Analysis
    Comparative Analysis: We use paired-samples t-tests to compare student performance on modular vs. project-based assessments and independent-samples t-tests to compare open-ended vs. interconnected projects. Effect sizes (Cohen’s d) are calculated.
@@
All analyses were conducted using Python with the SciPy and statsmodels libraries. Visualizations were generated using Matplotlib and Seaborn.
@@ IV Results
IV-A The AI Inflation Effect: Modular vs. Proctored Assessments
In the Spring 2025 Python for Data Science and AI course (N=41), we observed a stark contrast in student performance between modular assessments (knowledge checks and homework, where AI use was permitted) and proctored exams (where AI use was prohibited). As shown in Figure 1, students achieved near-perfect scores on knowledge checks (M=92.55, SD=20.26), but their performance dropped significantly on the proctored exams (M=62.82, SD=18.49). This represents a performance gap of nearly 30 percentage points, with a large effect size (Cohen’s d = 1.52).

### v3

@@ Abstract
The proliferation of generative AI tools has rendered traditional modular assessments in computing and data-centric education obsolete, creating a critical disconnect between academic training and industry practice. This paper presents a theoretically grounded framework for designing AI-resilient assessments, supported by proofs and empirical validation. We make three primary contributions. First, we establish two formal propositions: (1) assessments composed of interconnected problems, where outputs serve as inputs to subsequent stages, are inherently more AI-resilient than modular assessments due to multi-step reasoning and context limitations of large language models; and (2) semi-structured problems with deterministic success criteria provide more reliable measures of student competency than fully open-ended projects, which allow AI systems to default to familiar solution patterns. These findings challenge widely cited recommendations—discussed in a UNESCO-hosted analysis [39] and institutional guidelines [27, 14, 25]—that open-ended assessments maintain academic integrity and encourage deeper material engagement in the AI era—an assumption our findings contradict. Second, we validate these propositions through empirical analysis of three university data science courses ($N=117$), demonstrating a substantial AI inflation effect: students achieve near-perfect scores (9090100100%) on AI-assisted modular homework but drop approximately $30$ percentage points on proctored exams (Cohen’s $d=1.51$). Critically, interconnected projects maintain strong alignment with modular assessments ($r=0.954$, $p<.001$) while remaining AI-resilient, whereas proctored exams show only moderate correlation ($r=0.726$, $p<.001$), suggesting they assess different competencies. Third, we operationalize these findings into a practical design procedure that translates theoretical principles into implementable strategies, providing computing educators with a systematic, evidence-based approach for designing assessments that encourage deeper engagement, authentically reflect industry practice, and naturally resist trivial AI delegation.
@@ I Introduction
Since the release of generative AI tools like ChatGPT in late 2022 [28], the rapid adoption of this technology has reshaped professional practice at a fundamental level [33, 15]. Tasks that once defined early-career technical work—drafting boilerplate code, writing simple SQL queries, or executing small-scale data analysis—are now completed instantly by AI tools, a shift confirmed by large-scale industry experiments [11]. Yet, these same modular tasks continue to dominate academic assessments, creating a widening and unsustainable gap between university curricula and actual industry requirements [21]. This misalignment raises a critical question: What is the educational value of practicing skills that industry has already automated? Students quickly recognize this disconnect, undermining their motivation and learning when assignments feel obsolete [12].
@@
While educators widely acknowledge this gap, proposed solutions often focus on how instructors can use AI as a teaching assistant—for intelligent tutoring, adaptive feedback, or automated grading—rather than on how assessment design itself must fundamentally change [41, 5]. Such an approach overlooks the core challenge: the nature of meaningful technical work has shifted. Professionals now focus less on executing discrete tasks and more on orchestrating complex, multi-step workflows that integrate AI as a component, a form of higher-order work that current AI systems cannot replicate [7]. This shift demands a new pedagogical focus on AI-resilient assessments: assignments that cultivate genuine learning precisely because they cannot be fully delegated to an AI. Our approach to AI-resilient assessment design is grounded in recognized intrinsic limitations of large language models (LLMs) based on the transformer architecture. While proficient at single-step, modular problems, even state-of-the-art models like GPT-5 exhibit brittleness in tasks requiring sustained, multi-step reasoning [30, 32]. Empirical benchmarks consistently show that LLMs struggle with long inference chains and discovering abstract structural patterns, often exploiting superficial token biases rather than engaging in genuine inference [29, 16, 20]. These limitations are not merely theoretical; domain-specific evaluations in fields like biostatistics confirm that LLMs fail on multi-stage procedures unless guided by iterative human feedback [19]. Importantly, these constraints are architectural rather than incidental—rooted in how transformers process sequential information—and are unlikely to be resolved through scaling alone [30]. The boundary of current AI capability in multi-step and complex reasoning, therefore, defines a durable frontier for cultivating essential human skills that we must impart to our students.
@@
This paper argues that practical AI-resilient assessment design can be achieved by imbuing assessments with a high degree of interconnectedness among their components. Based on observations from teaching three sessions of upper-level courses in data science and statistical learning, we make two central claims. First, we contend that AI-resilient assessments composed of interconnected rather than modular problems foster integrative problem-solving skills that mirror industry workflows more closely—skills that current AI tools lack and that industry now demands. Second, we challenge widely cited recommendations discussed in a UNESCO-hosted analysis [39] and several institutional guidelines [27, 14, 25]—that open-ended assessments are a primary solution to maintaining academic integrity in the age of generative AI. While these recommendations rest on the assumption that AI "struggles with original thought and complex synthesis" [39, 27, 14, 25], recent evidence demonstrates that LLMs’ limitations stem not from an inability to generate creative responses, but from fundamental architectural constraints in multi-step reasoning and context management [30, 32].
@@ II Related Work
Our work is situated at the intersection of three distinct but related research areas: ($1$) the impact of generative AI on academic integrity and assessment, including recent evidence that generative AI can harm learning when used without appropriate guardrails [2, 9, 34, 40]; ($2$) the principles of authentic assessment and complex problem-solving in education [38, 10, 17]; and ($3$) the cognitive science of task complexity [8, 22, 3]. While no prior work directly proposes a framework for AI-resilience based on interconnectedness, these adjacent fields establish the foundation and necessity for our contribution.
@@ II-A Generative AI, Assessment, and Academic Integrity
The release of ChatGPT precipitated a surge of academic literature focused on the challenges generative AI poses to traditional assessment and academic integrity. A significant body of this work documents the scope of the problem, highlighting the ease with which students can use AI to complete assignments, thereby undermining learning and creating an integrity crisis [9, 34, 4]. These studies effectively diagnose the issue, arguing that the availability of powerful AI necessitates a fundamental rethinking of assessment practices. However, the proposed solutions often remain high-level, advocating for a shift towards in-class assessments, oral examinations, or a greater emphasis on process over product, without offering a systematic, scalable framework for redesigning coursework itself [40].
Moreover, an overreliance on in-class assessments may test a different skill altogether—test-taking ability rather than the intended learning outcomes [2]. Forgoing take-home or project-based assessments deprives students of opportunities to practice skills that closely mimic real-world workflows employed by industry professionals [38, 1]. We argue that designing AI-resilient assessments with appropriate guardrails is more beneficial, as it preserves authentic learning experiences while addressing integrity concerns.
@@ II-B Authentic Assessment and Complex Problem-Solving
Our approach builds on a long tradition of research into authentic assessment and complex problem-solving. Authentic assessment emphasizes mirroring the challenges and standards of professional practice, a principle well-established in engineering and computing education [38]. Similarly, the field of complex problem-solving (CPS) focuses on assessing cognitive processes in dynamic, non-routine situations, which aligns with our goal of moving beyond static, modular tasks [10, 17]. While philosophically aligned, this body of work predates the generative AI era and thus does not explicitly address how to design assessments that are resilient to AI delegation. It provides the pedagogical why—the need for real-world problem-solving—but not the operational how in a world where AI can solve many seemingly complex but isolated tasks.
@@ II-C Task Complexity and Cognitive Load Theory
To ground our concept of interconnectedness, we draw on Cognitive Load Theory (CLT) [35], particularly the concept of element interactivity. CLT posits that the intrinsic difficulty of a task is determined by the number of interacting elements that must be simultaneously processed in working memory [8]. A task with low element interactivity (e.g., solving a series of independent coding problems) can be broken down and solved piece by piece. In contrast, a task with high element interactivity (e.g., debugging a system where a change in one module has cascading effects on others) requires a holistic understanding of the entire system. We argue that current generative AI excels at tasks with low element interactivity but struggles with high-interactivity problems due to intrinsic multi-step reasoning limitations and context memory constraints inherent to the transformer architecture [30, 32]. To accomplish such tasks, students must engage meaningfully and expend cognitive effort, thereby achieving genuine learning without trivial AI delegation.
Building on the foundation of these three research areas, we now formalize these principles through two propositions that establish the theoretical basis for AI-resilient assessment design.
@@ III Methodology: A Framework for AI-Resilient Assessment
Our methodology is grounded in a theoretical framework that we first establish through two propositions. We then proceed to validate these propositions through a multi-year empirical study.
@@ III-A Theoretical Framework
We begin by formally stating and proving two propositions that underpin our assessment design principles.
Proposition 1.
@@
Proof of Proposition 1.
Let an assessment be a sequence of problems $P=(p_{1},p_{2},...,p_{n})$. In a modular design, each $p_{i}$ is independent. In an interconnected design, the solution to $p_{i+1}$ is contingent upon the solution of $p_{i}$. A generative AI model must therefore solve the entire sequence $(p_{1},...,p_{i})$ to correctly solve $p_{i}$. This requirement increases the computational and reasoning load on the model in two primary ways. First, it necessitates multi-step, sequential reasoning, a known weakness of current large language models (LLMs) where performance degrades as the number of reasoning steps increases [30, 32]. Second, the cumulative context required to solve later problems in the sequence can exceed the model’s effective context window, leading to information loss and degraded performance [26]. Because a modular design allows each problem to be solved in a separate, stateless context, it avoids these vulnerabilities. Thus, an interconnected design is inherently more resilient to trivial AI delegation and better reflects the complex, stateful workflows common in professional practice [6]. ∎
Proposition 2.
@@
Proof of Proposition 2.
When presented with a fully open-ended problem (e.g., "analyze a dataset of your choice"), the unconstrained solution space allows the model to select a problem instance within its training distribution, where memorized patterns can be exploited as shortcuts [13]. Research demonstrates that LLMs consistently follow the path of least resistance, preferring low-perplexity sequences that align with their training data and avoiding complex reasoning when simpler alternatives exist [36, 31, 18]. Consequently, students can trivially delegate open-ended tasks to AI, leveraging the model’s ability to produce outputs that exhibit the surface characteristics of original thought without genuine novel reasoning [23, 24]. This delegation is difficult to police, as AI detection tools are notoriously unreliable and easily evaded through simple paraphrasing techniques [Krishna2023, Erol2025]. A semi-structured assessment with deterministic success criteria eliminates this degree of freedom, forcing engagement with the specific problem at hand rather than a self-selected, AI-tractable variant.
A semi-structured assessment addresses this limitation by constraining the solution space through two mechanisms: (1) specifying a particular dataset or problem instance, and (2) defining deterministic success criteria (e.g., "achieve $90\%$ F1-score on this held-out test set"). These constraints prevent students from trivially delegating the task to AI, as the specified problem instance may lie outside the model’s training distribution, where memorized solutions and dataset artifacts cannot be exploited. The deterministic evaluation metric further ensures that students must optimize for correctness—an objective, task-specific criterion—rather than relying on AI-generated outputs that merely appear plausible. By narrowing the problem space to a specific, well-defined task, semi-structured assessments compel meaningful student engagement and skill practice, yielding a more reliable measure of competency while encouraging deeper material engagement [38].
@@ III-B Empirical Validation
To validate these propositions, we analyzed assessment data from three offerings of upper-level undergraduate and graduate courses in data science spanning multiple academic years. This section details the course context, the assessment designs, and the data analysis procedures.
@@ III-B1 Study Context and Population
The study was conducted across three course offerings at a large research university between Fall 2024 and Spring 2025:
@@
  • Predictive Analytics (Fall 2025): The same undergraduate course (N=42) but with a mix of modular assignments and a redesigned, interconnected final project.
III-B2 Analysis
For each course, we collected anonymized student grade data for all assessments. Our analysis includes:
@@ III-B2 Analysis
    Comparative Analysis: We use paired-samples t-tests to compare student performance on modular vs. interconnected problems. Effect sizes (Cohen’s d) are calculated. We also present empirical observations comparing fully open-ended and semi-open-ended assessment formats, along with a discussion of their relative effectiveness as part of our comparative analysis.
@@ IV Results
IV-A Descriptive Analysis: The AI Inflation Effect
For our descriptive analysis, our goal is to compare student performance across assessment types. Specifically, we examine whether there are performance gaps between modular take-home assessments, where generative AI tool use is permitted, and proctored assessments, where generative AI tool use is prohibited. The Spring 2025 Python for Data Science and AI course is the only course in which a proctored exam was offered; therefore, we use it as the reference for a fair comparison.
In the Spring 2025 Python for Data Science and AI course (N=41), we observed a stark contrast in student performance between take-home modular assessments (knowledge checks and homework, where AI use was permitted) and proctored exams (where AI use was prohibited). As shown in Figure 1, students achieved near-perfect scores on knowledge checks (mean [M] = 92.55, standard deviation [SD] = 20.26), but their performance dropped significantly on the proctored exams (M=62.82, SD=18.49). This represents a performance gap of nearly 30 percentage points, with a large effect size (Cohen’s d = 1.52).